Use the Method of Distribution Functions 2. (5 marks) Consider a random variable Y with density function 3 v 0 .else Find the probability density function of U 4- r2
Answer #1

2)

The cummulative distibution function od U is given by:

F(U)=P(U <u) = P(4- y? <u) = P(Y?> 4-u)

= 1- P(Y? <4-u) = 1- P(-14-u<Y<4-u)

= 1 - [F(V4 – u) - F(-V4 - 0) = 1+ F(-V4 – u) – F(V4 - u)

Now,

1-) (n – b1)s - (1-72: (n – 11-)5 = (),1 = (n)s

=[\frac{3(4-u)}{16}].\frac{-1}{2\sqrt{4-u}}(-1)-\frac{3(4-u)}{16}.\frac{1}{\sqrt{4-u}}(-1)

34 – 03 [ 16 ], 0 <u < 4

which is the PDF of U.